# GARCH copulas and GARCH-mimicking copulas

### v1

@@ 3.3 Computational examples
    It is well known that, for a covariance stationary GARCH(1,1) process ($\alpha_{1}+\beta_{1}<1$), the squared process $(X_{t}^{2})_{t\in\mathbb{Z}}$ behaves like an ARMA(1,1) process with autoregressive parameter $\alpha_{1}+\beta_{1}$ and moving-average parameter $-\beta_{1}$; see, for example, McNeil et al., (2015, Section 4.2.2). Thus the value of $\alpha_{1}+\beta_{1}$ controls the persistence, or the rate at which serial dependencies in the squared (or absolute value) process die away with lag. For a fixed value of $\alpha_{1}+\beta_{1}$ the magnitude of the serial dependency at lag 1 increases with $\alpha_{1}$, so that the ARCH(1) special case actually has the strongest dependence and the copula tends towards independence as $\alpha_{1}\to 0$. The parameters we have chosen are not necessarily typical of real financial data, where the ARCH effect $\alpha_{1}$ might be smaller and the GARCH effect $\beta_{1}$ larger, but have instead been selected to give a picture that has more evident dependence. The main point is to show that the behaviour of the GARCH(1,1) copula is qualitatively similar to that of the ARCH(1) copula.
  4. 4.
@@
    ARCH(1) model with Gaussian innovations and leverage (symmetric distribution). Here we use a volatility function $\sigma(x)$ satisfying $\sigma(x)^{2}=\alpha_{0}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}$ for $\gamma_{1}>0$ so that a negative return produces a larger value of the volatility than a positive return of the same magnitude. From now on, we concentrate on contour plots of the copula only, as shown in the top left panel of Figure 3. It can be seen that leverage removes radial symmetry and v-symmetry, but there is still h-symmetry (symmetry in the line $v=0.5$).
  5. 5.
@@
    Classical ARCH(1) model with skewed Student innovations (asymmetric distribution but symmetric volatility). This is shown in the top right panel of Figure 3 for an innovation distribution that is skewed to the left (towards negative shocks). The skewness removes all the previous symmetries in the pictures.
  6. 6.
    Absolute value copula for classical ARCH(1) model with Gaussian innovations. This is shown in the bottom left of Figure 3 for the same process as in the top left of Figure 2. It can be thought of as a blown-up version of the top right quadrant of the latter and clearly shows non-exchangeability and stronger dependence in the upper tail. Note that this picture also illustrates the density of the volatility copula of ARCH(1), that is the copula of $(\sigma_{1},\sigma_{2})^{\top}$; this follows because $\sigma_{t}$ is a strictly increasing function of $|X_{t-1}|$ for all $t$ and copulas are invariant under strictly increasing transformations of variables. For the same reason, it is also the copula density for the variables $(X_{1}^{2},X_{2}^{2})^{\top}$.
@@
Figure 3: Introducing asymmetry. Top left: ARCH(1) with Gaussian innovations in which the volatility function has leverage and the parameters are $\alpha_{0}=0.4$, $\alpha_{1}=0.3$ and $\gamma_{1}=0.4$. Top right: classical ARCH(1) model with with 4 degrees of freedom and skewness parameter $\lambda=0.8$ where $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$. Bottom left: copula density for copula of the absolute values for classical ARCH(1) copula with Gaussian innovations. Bottom right: AR(1)-ARCH(1) model with Gaussian innovations and parameters $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
3.4 AR(1) model with ARCH(1) errors
We consider briefly a time series model with stochastic volatility as well as a non-zero stochastic conditional mean term. We consider an AR(1) model with ARCH(1) errors and show the implications of the conditional mean term. Specifically, we consider the first-order Markov process $(X_{t})_{t\in\mathbb{Z}}$ satisfying equations
$X_{t}=\phi X_{t-1}+\sigma(X_{t-1})\;\epsilon_{t}\quad\text{with}\quad\sigma(x)=\sqrt{\alpha_{0}+\alpha_{1}x^{2}},$ (11)
where the innovations $(\epsilon_{t})_{t\in\mathbb{Z}}$ have a symmetric distribution. This model was proposed and analysed by Borkovec and Klüppelberg, (2001) [^2] It differs from the usual version of the AR(1)-ARCH(1) model, $X_{t}=\phi X_{t-1}+\sigma(X_{t-1}-\phi X_{t-2})\;\epsilon_{t}$, but this is not first-order Markov and thus less straightforward to treat with our methods. who showed that it is covariance stationary provided $\phi^{2}+\alpha_{1}<1$ and has a symmetric stationary distribution $F_{X}$.
It may be easily verified that the copula density satisfies $c(u,v)=c(1-u,1-v)$ and so the copula is radially symmetric. However, the copula is neither h-symmetric nor v-symmetric, and thus it is not jointly symmetric. The copula density is shown in the bottom right panel of Figure 3 in the case where $\phi=0.3$, $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$. It is evident that more of the probability mass is now on the main diagonal of the copula.
In the remainder of the paper we concentrate on models without a conditional mean term, but this example gives some clues about the effect on the copula of serial dependence in the untransformed process $(X_{t})_{t\in\mathbb{Z}}$ in combination with stochastic volatility, a point we return to in our discussion in Section 7.
3.5 Higher-order conditional copulas of GARCH(1,1)-type processes
For higher-order copulas we introduce the compact notation $\bm{X}_{[j,l]}=(X_{j},\ldots,X_{l})^{\top}$ to denote random vectors consisting of a contiguous block of random variables; we apply the same notation to realisations of the random vectors. Whenever the notation is used with $j>l$, this means that the variables are taken in reverse order to their natural temporal order.
@@ 3.5 Higher-order conditional copulas of GARCH(1,1)-type processes
Now we consider the copula $C_{k}$ of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $\bm{X}_{[2:k]}=\bm{x}_{[2:k]}$ for $k\geqslant 2$. The density of the conditional copula takes the form
$c_{k}(u,v\mid\bm{x}_{[2:k]})=\frac{f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right),F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}{f_{X_{1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)f_{X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}$
@@
Note that this copula density will simply be the independence copula density for a process of ARCH(1) type, since $X_{1}$ and $X_{3}$ are conditionally independent given $X_{2}$. Consider, therefore a process with an explicit GARCH effect, i.e. $\sigma(x,s)$ depends also on $s$.
To develop intuition we consider the case where $k=2$. The symmetry properties can be investigated by considering the joint conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$. To this end consider the joint density
$f_{X_{1},X_{2},X_{3}}(x,z,y)$ $=\int_{0}^{\infty}f_{X_{1},X_{2},X_{3},\sigma_{1}}(x,z,y,s)\mathrm{d}s$
$=\int_{0}^{\infty}f_{X_{3}\mid X_{2},X_{1},\sigma_{1}}(y\mid z,x,s)f_{X_{2}\mid X_{1},\sigma_{1}}(z\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
$=\int_{0}^{\infty}\frac{1}{\sigma(z,\sigma(x,s))}f_{\epsilon}\left(\frac{y}{\sigma(z,\sigma(x,s))}\right)\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{z}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s.$
It is straightforward to see that a process with symmetric innovations will satisfy $f_{X_{1},X_{2},X_{3}}(x,z,-y)=f_{X_{1},X_{3},X_{2}}(x,z,y)$ while a symmetric process will also satisfy $f_{X_{1},X_{2},X_{3}}(-x,z,y)=f_{X_{1},X_{2},X_{3}}(x,z,y)$. These properties will also hold for the conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)=f_{X_{1},X_{2},X_{3}}(x,z,y)/f_{X}(z)$ and thus, by Lemma 1, the various forms of symmetry will be inherited by the conditional copula density $c_{2}(u,v\mid z)$. By applying a similar analysis to the joint density $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)$ we can therefore establish the following result.
@@
    In a GARCH(1,1)-type process with symmetric distribution, the conditional copula $C_{k}$ is h-symmetric about zero for $k\geqslant 2$.
@@
    In a symmetric GARCH(1,1)-type process the conditional copula is jointly symmetric about the origin for $k\geqslant 2$.
Proof.
Suppose we write $s_{2}(x,s)=\sigma(x,s)$, $s_{3}(z_{1},x,s)=\sigma(z_{1},s_{2}(x,s))$ and then for $j=4,\ldots,k+1$ we define recursive functions $s_{j}(z_{j-2},\ldots,z_{1},x,s)=\sigma(z_{j-2},s_{j-1}(z_{j-3},\ldots,z_{1},x,s))$. This permits us to write the joint density as
$f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)=\int_{0}^{\infty}f_{X_{k+1}\mid\sigma_{k+1}}(y\mid s_{k+1}(z_{k-1},\ldots,z_{1},x,s))\\ \left(\prod_{j=3}^{k}f_{X_{j}\mid\sigma_{j}}(z_{j-1}\mid s_{j}(z_{j-2},\ldots,z_{1},x,s))\right)f_{X_{2}\mid\sigma_{2}}(z_{1}\mid s_{2}(x,s))f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
It is clear that this joint density satisfies $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(x,\ldots,-y)$ when the innovation distribution is symmetric. Moreover it is also clear that in the symmetric model we have $f_{X_{1},\ldots,X_{k+1}}(x,\ldots,y)=f_{X_{1},\ldots,X_{k+1}}(-x,\ldots,y)$ since all of the functions $s_{2},\ldots,s_{k+1}$ are even functions of $x$. These properties will also hold for the conditional density
$f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}(x,y\mid z_{1},\ldots,z_{k-1})=f_{X_{1},\ldots,X_{k+1}}(x,z_{1},\ldots,z_{k-1},y)/f_{X_{2},\ldots,X_{k}}(z_{1},\ldots,z_{k-1})$
and thus we can apply Lemma 1 to infer that the assertions regarding symmetry of the copula hold. ∎
Remark 1.
Since the joint density of $(X_{1},X_{k+1})^{\top}$ can be obtained as
it follows that the same method of proof can be used to show that the copula of the unconditional distribution of $(X_{1},X_{k+1})^{\top}$ shares the same symmetry properties as the conditional copula.
An interesting question concerns the degree to which $c_{k}(u,v\mid\cdot)$ depends on the realized values of the conditioning variables. We investigate this question for $k=2$ and and write $c_{2}(u,v;w):=c_{2}(u,v\mid F_{X}^{-1}(w))$. The latter function takes the form
$c_{2}(u,v;w)=\frac{f_{X_{1},X_{3}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right),F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}{f_{X_{1}\mid X_{2}}\left(F_{X_{1}\mid X_{2}}^{-1}\left(u\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)f_{X_{3}\mid X_{2}}\left(F_{X_{3}\mid X_{2}}^{-1}\left(v\mid F_{X}^{-1}(w)\right)\mid F_{X}^{-1}(w)\right)}$
and can be calculated approximately for the classical GARCH(1,1) model building on our analysis in Section 3.3.
In Figure 4 we show contour plots for the cases where $w=0.5$, i.e. the intermediate observation is at the median of the marginal distribution (zero in this case), and $w=0.75$, i.e. the intermediate observation is at a quartile of the marginal distribution. While the two pictures are not identical, they are relatively similar to each other in terms of the shapes and heights of the contour lines. To examine this further we attempt to compute an approximate value for the absolute distance between the conditional copula densities and the independence copula density using the measure $D_{2}(w)=\int_{0}^{1}\int_{0}^{1}|c_{2}(u,v;w)-1|\mathrm{d}u\mathrm{d}v$. We obtain the estimates $D_{2}(0.5)=0.10$ and $D_{2}(0.75)=0.09$, which can be compared with the equivalent measure $D_{1}=0.16$ for the copula density $c_{1}$ in Figure 2.
As might be expected, the dependence weakens from lag 1 to lag 2. Moreover, the value of the conditioning variable does not seem to have a particularly large effect on the conditional copula at lag 2. This might be taken as weak evidence that the simplifying assumption is approximately tenable; obviously we should look at values of $w$ further into the tail, but our approximation methods then become more unreliable.
Figure 4: Contour plots of $c_{2}(u,v;w)$ for the classical GARCH(1,1) process with Gaussian innovations. In the left-hand figure $w=0.5$ and in the right-hand figure $w=0.75$.
4 Copula densities and v-transforms
@@ 4.1 V-transforms and their inverses
The function $\mathcal{V}(u)=|2u-1|$ which appeared in Proposition 3 is a symmetric v-shaped function and the simplest example of a v-transform. These are special cases of uniform-measure-preserving transformations (Porubský et al.,, 1988) and are extensively analysed in McNeil, (2021) and Bladt and McNeil, 2022b . In this section we recap the essential elements of the theory and examine its relevance for understanding and mimicking the behaviour of the copula of processes of GARCH(1,1) type.
@@
for a parameter $10<\delta<$ 1, known as the fulcrum, and a continuous and strictly increasing distribution function $\Psi$ on $[0,1]$, known as the generator.
Mappings of this form are always v-shaped with $\mathcal{V}(0)=\mathcal{V}(1)=1$ and $\mathcal{V}(\delta)=0$. They preserve uniformity in the sense that if $U\sim U(0,1)$, then $\mathcal{V}(U)\sim U(0,1)$. If the generator $\Psi$ is taken to be standard uniform, we obtain the linear v-transform, which satisfies $\mathcal{V}(u;\delta)=(\delta-u)/\delta$ for $u\leqslant\delta$ and $\mathcal{V}(u;\delta)=(u-\delta)/(1-\delta)$ for $u>\delta$; this includes the special case $\mathcal{V}(u;0.5)=|2u-1|$, which is the symmetric v-transform. It is possible to introduce further parameters. For example, by taking $\Psi(x)=x^{\kappa}$ for $\kappa>0$, we obtain a two-parameter family
@@
For further results we need some more notation. For every $y\in(0,1]$ there are two possible values $u$ on either side of the fulcrum $\delta$ such that $\mathcal{V}(u)=y$ and these may be written as $\mathcal{V}^{-1}(y)$ and $\mathcal{V}^{-1}(y)+y$ where $\mathcal{V}^{-1}(y)=\inf\{u:\mathcal{V}(u)=y\}$ is a partial inverse of $\mathcal{V}$. We also introduce the function $\Delta(y)=-1/\mathcal{V}^{\prime}(\mathcal{V}^{-1}(y))$ where $\mathcal{V}^{\prime}$ is the derivative of $\mathcal{V}$. If $U\sim U(0,1)$ then McNeil, (2021) shows that $\Delta(y)=\mathbb{P}(U\leqslant\delta\mid\mathcal{V}(U)=y)$ for $y\in(0,1]$; for the linear v-transform $\Delta(y)=\delta$, but otherwise $\Delta(y)$ depends explicitly on $y$.
We can stochastically invert a v-transform in such a way that we also preserve uniformity of random variables. For $y\in(0,1]$, the stochastic inverse $\mathcal{V}^{\leftarrow}(y)$ of the v-transform $\mathcal{V}$ is defined to take the value $\mathcal{V}^{-1}(y)$ (from the left branch of the transform) with probability $\Delta(y)$ and the value $\mathcal{V}^{-1}(y)+y$ (from the right branch of the transform) with probability $1-\Delta(y)$; for $y=0$ we set $\mathcal{V}^{\leftarrow}(y)=\delta$. If $U\sim U(0,1)$ then it can also be shown that $\mathcal{V}^{\leftarrow}(U)\sim U(0,1)$. Note, however, that while $\mathcal{V}\circ\mathcal{V}^{\leftarrow}(y)=y$ holds for all $y\in[0,1]$, the equation $\mathcal{V}^{\leftarrow}\circ\mathcal{V}(u)=u$ is only true with probability $\Delta(u)$, if $u<\delta$, and with probability $1-\Delta(u)$, if $u>\delta$.

### v2

@@ 2.4.3 Classical GARCH(1,1) process with Gaussian innovations (symmetric)
This is shown in the bottom right panel of Figure 1 for a process with Gaussian innovations. It is well known that, for a covariance stationary GARCH(1,1) process ($\alpha_{1}+\beta_{1}<1$), the squared process $(X_{t}^{2})_{t\in\mathbb{Z}}$ behaves like an ARMA(1,1) process with autoregressive parameter $\alpha_{1}+\beta_{1}$ and moving-average parameter $\beta_{1}$; see, for example, McNeil et al., (2015, Section 4.2.2). Thus the value of $\alpha_{1}+\beta_{1}$ controls the persistence, or the rate at which serial dependencies in the squared (or absolute value) process die away with lag. For a fixed value of $\alpha_{1}+\beta_{1}$ the magnitude of the serial dependency at lag 1 increases with $\alpha_{1}$, so that the ARCH(1) special case actually has the strongest dependence and the copula tends towards independence as $\alpha_{1}\to 0$. The parameters we have chosen are not necessarily typical of real financial data, where the ARCH effect $\alpha_{1}$ might be smaller and the GARCH effect $\beta_{1}$ larger, but have instead been selected to give a picture that has more evident dependence. The main point is to show that the behaviour of the GARCH(1,1) copula is qualitatively similar to that of the ARCH(1) copula.
2.4.4 ARCH(1) model with Gaussian innovations and leverage (symmetric distribution)
@@
In the left panel of Figure 2 we show the copula $c_{1}$ for a model with symmetric innovations and volatility function $\sigma(x)$ satisfying $\sigma(x)^{2}=\alpha_{0}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}$ for $\gamma_{1}>0$, so that a negative return produces a larger value of the volatility than a positive return of the same magnitude. It can be seen that leverage removes radial symmetry and v-symmetry, but there is still h-symmetry (symmetry in the line $v=0.5$).
2.4.5 Classical ARCH(1) model with skewed Student innovations (symmetric volatility)
@@
The copula density $c_{1}$ is shown in the right panel of Figure 2 for a model with an innovation distribution that is skewed to the left (towards negative shocks). The skewness removes all the previous symmetries in the pictures.
@@
Figure 2: Introducing asymmetry. Left: ARCH(1) with Gaussian innovations in which the volatility function has leverage and the parameters are $\alpha_{0}=0.4$, $\alpha_{1}=0.3$ and $\gamma_{1}=0.4$. Right: classical ARCH(1) model with 4 degrees of freedom and skewness parameter $\lambda=0.8$ where $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$.
2.5 Higher-order copulas of GARCH(1,1)-type processes
@@
In this section we show that the same symmetries evident in the copula density $c_{1}$ can be found in the density $c_{k}$ of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $\bm{X}_{[2:k]}=\bm{x}_{[2:k]}$ for $k\geqslant 2$, which takes the form
$c_{k}(u,v\mid\bm{x}_{[2:k]})=\frac{f_{X_{1},X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right),F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}{f_{X_{1}\mid\bm{X}_{[2:k]}}\left(F_{X_{1}\mid\bm{X}_{[2:k]}}^{-1}\left(u\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)f_{X_{k+1}\mid\bm{X}_{[2:k]}}\left(F_{X_{k+1}\mid\bm{X}_{[2:k]}}^{-1}\left(v\mid\bm{x}_{[2:k]}\right)\mid\bm{x}_{[2:k]}\right)}$ (10)
It should be distinguished from the copula density $c_{(k)}$ of the copula $C_{(k)}$ of the marginal distribution of $(X_{1},X_{k+1})^{\top}$ (unless $k=1$). In particular, it is the densities $c_{k}$ rather than $c_{(k)}$ that appear in the D-vine representation (1) of the GARCH-type process. The former depend on the conditioning variables $\bm{x}_{[2:k]}$ although we would like to be able to assume that this dependence is weak so that the D-vine representation can be approximated by a simplified D-vine.
@@
To understand the main argument in this section it suffices to consider the case $k=2$. We consider a process with an explicit GARCH effect since, for a process of ARCH(1) type, $X_{1}$ and $X_{3}$ are conditionally independent given $X_{2}=z$ and $c_{2}(u,v\mid z)\equiv 1$, the density of the independence copula. The joint conditional density $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$ and the joint density $f_{X_{1},X_{3}}(x,y)$ can both be calculated from the joint density $f_{X_{1},X_{2},X_{3}}(x,z,y)$ by observing that $f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)=f_{X_{1},X_{2},X_{3}}(x,z,y)/f_{X}(z)$ and $f_{X_{1},X_{3}}(x,y)=\int_{-\infty}^{\infty}f_{X_{1},X_{2},X_{3}}(x,z,y)\,\mathrm{d}z$. Since
$f_{X_{1},X_{2},X_{3}}(x,z,y)$ $=\int_{0}^{\infty}f_{X_{3}\mid X_{2},X_{1},\sigma_{1}}(y\mid z,x,s)f_{X_{2}\mid X_{1},\sigma_{1}}(z\mid x,s)f_{X_{1}\mid\sigma_{1}}(x\mid s)f_{\sigma}(s)\mathrm{d}s$
$=\int_{0}^{\infty}\frac{1}{\sigma(z,\sigma(x,s))}f_{\epsilon}\left(\frac{y}{\sigma(z,\sigma(x,s))}\right)\frac{1}{\sigma(x,s)}f_{\epsilon}\left(\frac{z}{\sigma(x,s)}\right)\frac{1}{s}f_{\epsilon}\left(\frac{x}{s}\right)f_{\sigma}(s)\mathrm{d}s$
we can infer that $f_{X_{1},X_{2},X_{3}}(x,z,-y)=f_{X_{1},X_{3},X_{2}}(x,z,y)$ for a process with symmetric innovations and $f_{X_{1},X_{2},X_{3}}(-x,z,y)=f_{X_{1},X_{2},X_{3}}(x,z,y)$ for a symmetric process. It follows immediately that the conditional and unconditional densities satisfy $f_{X_{1},X_{3}\mid X_{2}}(x,-y\mid z)=f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$ and $f_{X_{1},X_{3}}(x,-y)=f_{X_{1},X_{3}}(x,y)$ for a process with symmetric innovations and we have h-symmetry in both cases. Moreover, they satisfy $f_{X_{1},X_{3}\mid X_{2}}(-x,y\mid z)=f_{X_{1},X_{3}\mid X_{2}}(x,y\mid z)$ and $f_{X_{1},X_{3}}(-x,y)=f_{X_{1},X_{3}}(x,y)$ for a symmetric process and we have joint symmetry in both cases. By Lemma 1, the various forms of symmetry are inherited by the conditional copula $C_{2}$ of $(X_{1},X_{3})^{\top}$ given $X_{2}$ and the marginal copula $C_{(2)}$ of $(X_{1},X_{3})^{\top}$.
We now give the general result for $k\geqslant 2$. Since the logic of its proof is identical to the argument given above, albeit with more cumbersome notation, the general proof is given in the Supplementary Material.
@@
    In a GARCH(1,1)-type process with symmetric distribution, the conditional copula $C_{k}$ and the marginal copula $C_{(k)}$ are h-symmetric about zero for $k\geqslant 2$.
@@
    In a symmetric GARCH(1,1)-type process $C_{k}$ and $C_{(k)}$ are jointly symmetric about the origin for $k\geqslant 2$.
The computational tools we have developed to approximate $c_{1}$ in Section 2.4 can be extended to approximate $c_{2}$, and we use them to explore the simplifying assumption. Let $c_{2}(u,v;w):=c_{2}(u,v\mid F_{X}^{-1}(w))$ for $w\in(0,1)$. Contour plots of $c_{1}(u,v;w)$ for the cases where $w=0.5$ and $w=0.75$ are shown in Figure 3. The two pictures are very similar to each other in terms of the shapes and heights of the contour lines, and are also similar in pattern to the plot of $c_{1}$ in the lower-right panel of Figure 1. We compute an approximate value for the absolute distance between the conditional copula densities and the independence copula density using the measure $D_{2}(w)=\int_{0}^{1}\int_{0}^{1}|c_{2}(u,v;w)-1|\mathrm{d}u\mathrm{d}v$. We obtain the estimates $D_{2}(0.5)=0.10$ and $D_{2}(0.75)=0.09$, which can be compared with the equivalent measure $D_{1}=0.16$ for the copula density $c_{1}$ in the lower-right panel of Figure 1.
Figure 3: Contour plots of $c_{2}(u,v;w)$ for the classical GARCH(1,1) process with Gaussian innovations. Left: $w=0.5$. Right: $w=0.75$.
As might be expected, the dependence weakens from lag 1 to lag 2. Moreover, the value of $w$ does not seem to have a particularly large effect on the conditional copula at lag 2. This might be taken as weak evidence in favour of the simplifying assumption. It would be desirable to look at values of $w$ further into the tail and consider lags $k>2$, but our approximation methods then become less accurate leading to less confident conclusions.
3 Using v-transforms to approximate GARCH pair copulas
@@ 3.1 V-transforms and their inverses
It has long been observed that serial dependence in financial asset-price return series, and in the models like GARCH that describe them, shows up as serial correlation of absolute or squared values. Loaiza-Maya et al., (2018) considered the componentwise application of symmetric v-shaped transformations $T$ (such as the absolute value function $T(x)=|x|$) to a pair of random variables $(Y,Z)^{\top}$ and derived an expression for the relationship of the copula $C_{T}$ of $(T(Y),T(Z))^{\top}$ to the copula $C$ of $(Y,Z)^{\top}$. They showed that the resulting copula $C_{T}$ involves the marginal distributions $F_{Y}$ and $F_{Z}$ except in the special case of symmetric margins. McNeil, (2021) observed that the effect on the copula of such transformations, as well as more general asymmetric transformations, can be isolated by introducing v-shaped transformations of uniform random variables that are uniform-distribution-preserving (Porubský et al.,, 1988). Importantly, for the purposes of this paper, these so-called v-transforms can be inverted stochastically in a way that also preserves uniform distributions. We use this inversion technique to create new copulas that show behaviour similar to GARCH pair copulas. In this section we recap the essential elements of the theory.
@@
for a parameter $0<\delta<1$, known as the fulcrum, and a continuous and strictly increasing distribution function $\Psi$ on $[0,1]$, known as the generator.
Mappings of this form are always v-shaped with $\mathcal{V}(0)=\mathcal{V}(1)=1$ and $\mathcal{V}(\delta)=0$. They preserve uniformity in the sense that if $U\sim\mathcal{U}(0,1)$ is standard uniform, then $\mathcal{V}(U)\sim\mathcal{U}(0,1)$. If the generator is $\Psi(x)=x$ (the uniform df), we obtain the linear v-transform, which satisfies $\mathcal{V}(u;\delta)=(\delta-u)/\delta$ for $u\leqslant\delta$ and $\mathcal{V}(u;\delta)=(u-\delta)/(1-\delta)$ for $u>\delta$; this includes the special case $\mathcal{V}(u;0.5)=|2u-1|$, which is the symmetric v-transform. It is possible to introduce further parameters. For example, by taking $\Psi(x)=x^{\kappa}$ for $\kappa>0$, we obtain a two-parameter family
@@
For further results we need some more notation. For every $y\in(0,1]$ there are two possible values $u$ on either side of the fulcrum $\delta$ such that $\mathcal{V}(u)=y$ and these may be written as $\mathcal{V}^{-1}(y)$ and $\mathcal{V}^{-1}(y)+y$ where $\mathcal{V}^{-1}(y)=\inf\{u:\mathcal{V}(u)=y\}$ is a partial inverse of $\mathcal{V}$. We also introduce the function $\Delta(y)=-1/\mathcal{V}^{\prime}(\mathcal{V}^{-1}(y))$ where $\mathcal{V}^{\prime}$ is the derivative of $\mathcal{V}$. If $U\sim\mathcal{U}(0,1)$ then McNeil, (2021) shows that $\mathbb{P}(U\leqslant\delta\mid\mathcal{V}(U)=y)=\Delta(y)$ for $y\in(0,1]$; for the linear v-transform $\Delta(y)=\delta$, but otherwise $\Delta(y)$ depends explicitly on $y$.
We can stochastically invert a v-transform in such a way that we also preserve uniformity of random variables. For $y\in(0,1]$, the stochastic inverse $\mathcal{V}^{\leftarrow}(y,W)$ of the v-transform $\mathcal{V}$ is defined by taking a uniform random variable $W\sim\mathcal{U}(0,1)$ and setting $\mathcal{V}^{\leftarrow}(y,W)=\mathcal{V}^{-1}(y)$ if $W\leqslant\Delta(y)$ and $\mathcal{V}^{\leftarrow}(y,W)=\mathcal{V}^{-1}(y)+y$ otherwise; we always set $\mathcal{V}^{\leftarrow}(0,W)=\mathcal{V}^{-1}(0)=\delta$. If $U$ and $W$ are iid $\mathcal{U}(0,1)$ variables, then it can also be shown that $\mathcal{V}^{\leftarrow}(U,W)\sim\mathcal{U}(0,1)$. Note, however, that while $\mathcal{V}\left(\mathcal{V}^{\leftarrow}(y,W)\right)=y$ holds for all $y\in[0,1]$, the equation $\mathcal{V}^{\leftarrow}\left(\mathcal{V}(u),W\right)=u$ is only true with probability $\Delta(u)$, if $u<\delta$, and with probability $1-\Delta(u)$, if $u>\delta$.